# Zens, Díaz, Durante and Patacchini

### v1

@@ Abstract
Criminal activity data are typically available via a three-way tensor encoding the reported frequencies of different crime categories across time and space. The challenges that arise in the design of interpretable, yet realistic, model-based representations of the complex dependencies within and across these three dimensions have led to an increasing adoption of black-box predictive strategies. Although this perspective has proved successful in producing accurate forecasts guiding targeted interventions, the lack of interpretable model-based characterizations of the dependence structures underlying criminal activity tensors prevents from inferring the cascading effects of these interventions across the different dimensions. We address this gap through the design of a low-rank bilinear autoregressive model which achieves comparable predictive performance to black-box strategies, while allowing interpretable inference on the dependence structures of criminal activity reports across crime categories, time, and space. This representation incorporates the time dimension via an autoregressive construction, accounting for spatial effects and dependencies among crime categories through a separable low-rank bilinear formulation. When applied to Chicago police reports, the proposed model showcases remarkable predictive performance and also reveals interpretable dependence structures unveiling fundamental crime dynamics. These results facilitate the design of more refined intervention policies informed by cascading effects of the policy itself.
@@
The empirical analysis of criminal activities plays an important role in both theory development and public policy considerations. Theoretical models of crime indicate that criminal incidents do not occur randomly, but rather exhibit complex interdependencies across spatial areas, time, and crime categories (e.g., Brantingham and Brantingham, 2016; Weisburd, 2015). As a result, several methods have been proposed to characterize the intricate system of interdependencies within and across these three dimensions. Available solutions for addressing this endeavor range from interpretable model-based representations (e.g., Shoesmith, 2013; Law et al., 2014; Liu and Zhu, 2017; Quick et al., 2018; Clark and Dixon, 2018; Balocchi and Jensen, 2019; Vicente et al., 2023; Balocchi et al., 2023) to more sophisticated black-box predictive strategies (e.g., Guo and Wu, 2012; Chun et al., 2019; Walczak, 2021; Berk, 2021; Zhang et al., 2022; Liang et al., 2022; Mandalapu et al., 2023; Wu et al., 2024; Wang et al., 2025). Although statistical models allow, in principle, for interpretable inference on the complex dependence structures underlying the observed criminal activity tensors, the validity of the final conclusions is often undermined by the challenges that arise in the design of sufficiently-flexible, yet interpretable, realistic model-based representations for these dependence structures. As such, the existing statistical models often constrain the analysis by focusing on two dimensions of the three-way tensor, fixing the third at a selected crime category, a single geographic location, or a pre-specified time interval. Although such a perspective has led to important advancements, as discussed above, constraining the analysis only to two dimensions fails to characterize the full set of interdependencies underlying the criminal activity tensor, thereby limiting the policy impact of the resulting inferences. Addressing these issues would require extending the available flexible models for two dimensions to similarly-effective formulations for the entire three-way tensor. In fact, instead of exploring such a direction, most efforts in the recent years have been devoted to the development and implementation of black-box strategies within a predominantly prediction-oriented framework (see, e.g., Berk, 2021; Mandalapu et al., 2023). These strategies have proved successful in producing accurate forecasts guiding targeted intervention policies, such as hot spot policing (Weisburd and Telep, 2014; Braga et al., 2014), but generally lack a model-based construction, and therefore make it challenging to derive meaningful structural insights on the cascading, positive or negative, spillover effects of specific localized interventions (see, e.g., Weisburd et al., 2006; Blattman et al., 2021). Learning these effects is fundamental to devise improved intervention policies that not only rely on the forecasted criminal activities, but are also guided by the expected cascading effects of the planned policy itself.
@@
Motivated by the above discussion and challenges, we design within Section 3 a model-based Bayesian representation of three-way criminal activity tensors that achieves a predictive performance comparable to routinely-used black-box strategies, while allowing for interpretable inference on the complex dependence structures within and across crime categories, spatial locations and time intervals. This formulation is guided by qualitative and quantitative theories of crime discussed in Section 2, and relies on a temporal autoregressive construction that incorporates spatial effects and graphical dependencies among different crime categories via a separable low-rank bilinear formulation inspired by the general class of multilinear tensor regression models in Hoff (2015). Besides reducing the dimensionality of the problem, this novel perspective to the analysis of crime activity tensors crucially unveils sparse averaged inter-temporal dependencies between crime categories within spatial locations as well as among spatial locations for each crime category. Such a separable structure is also considered for the noise component of the model, thereby further uncovering across-categories and across-locations immediate reactions to shocks, possibly arising from targeted hot spot interventions. As clarified in Section 4, inference under the proposed model is performed through a Bayesian approach leveraging a suitably-designed Gibbs sampling algorithm. The advantages of our formulation are illustrated within Section 5 in an application to a three-way criminal activity tensor obtained from the Chicago Police Department. Besides achieving a predictive performance comparable to routinely-implemented black-box strategies, the proposed model unveils yet-unexplored statistical interdependencies between crime categories and within locations over time that augment the set of information available to those policymakers seeking to understand and prevent complex patterns in criminal activities. As discussed in the concluding remarks within Section 6, our novel approach can be applied in general to criminal activity tensors from any urban environment for which similar data are available.
@@
Crime is mainly an urban phenomenon. Urban agglomerations typically exhibit higher crime rates compared to smaller towns and rural areas (e.g., Glaeser and Sacerdote, 1999). Although larger cities are often associated with higher economic development, and hence more attractive, this is not the only dimension that explains the distribution of crime across cities. In fact, crime reports are typically sorted unequally across different urban locations, with offenses often concentrated by type in relatively small areas. This spatial heterogeneity, coupled with specific temporal persistences, has led social scientists to focus on local determinants rather than individual drivers.
Three traditional criminology perspectives are useful for understanding the connections among time, space and crime categories. Crime opportunity theories, which focus on the factors shaping space-time crime patterns, particularly the clustering of criminal activities (e.g., Brantingham and Brantingham, 2017; Cohen and Felson, 1979); social control theories, where the structural dimensions of a neighborhood influence residents propensity to exert informal social control (Sampson and Groves, 1989; Shaw and McKay, 1942); and theories that simply highlight how local characteristics (e.g., medical marijuana dispensaries, parks, neighborhood stability and socioeconomic disparities) affect incentives for committing a specific crime within a given area (Contreras, 2017; Contreras and Hipp, 2020; Kubrin and Hipp, 2016). It is important to note that crime clusters are not driven by a single factor in isolation (Hipp and Williams, 2020). Rather, the interplay between opportunity, social controls, and local conditions causes urban areas to experience varying levels of temporal exposure to specific types of crimes and violence.
@@
Although it is challenging to separate time and space when analyzing urban crime patterns, the temporal dimension of crime determinants has been documented to play a crucial role in shaping its spatial distribution, alongside factors related to the offender behavior. Several criminogenic features of a geographic area tend to remain stable over time, contributing to the persistence of specific criminal activities (Johnson, 2008). For example, neighborhood characteristics that either promote or inhibit social control are likely to remain constant (Sampson and Groves, 1989). In addition, the success of previous offenses often motivates criminals to continue operating in the same area, capitalizing on the knowledge acquired about available targets and absence of countermeasures (Cohen and Felson, 1979). The literature on near repeat victimization further suggests that offenders are more likely to target locations where similar crimes have been successfully committed in the past (Johnson and Bowers, 2004; Townsley et al., 2003). However, not all crimes are equally susceptible to temporal persistence and spatial clustering (e.g., Pease et al., 1998; Johnson and Bowers, 2004; Ratcliffe and Rengert, 2008; Johnson et al., 2009; Haberman and Ratcliffe, 2012; Amemiya et al., 2020; Chainey, 2021; Chainey and Muggah, 2022). In fact, as discussed in these studies, the clustering of urban crime in time and space is shaped by patterns that are heterogeneous across crime categories. In specific urban areas, certain types of offenses are more prevalent than others, leading to a localized concentration of particular crimes (e.g., property crimes versus violent crimes). These patterns are shaped by the spatial distribution of suitable targets and by the presence of deterrence interventions (Cohen and Felson, 1979), which can favor the occurrence of some crimes over others. For instance, the covid-19 pandemic highlighted the role of crime opportunity in explaining the global reduction in street crime (Nivette et al., 2021; Trajtenberg et al., 2024), while at the same time, the literature documented increases in domestic violence and cybercrime (Piquero et al., 2021; Ravindran and Shah, 2023). Empirical evidence suggests that offenders may have increased the use of violence to compensate for the reduced opportunities brought on by the pandemic, thereby reshaping the spatial distribution and clustering of violent and non-violent crimes (Díaz et al., 2022). Another example of crime-type clustering heterogeneity results from the strategic decision-making of organized crime groups when selecting or relocating to new territories (Varese, 2011). These groups typically target urban communities that offer favorable conditions for the development of illegal markets, display socio-economic deprivation, and demonstrate a higher degree of homophily (Campana and Meneghini, 2024; Katz and Schnebly, 2011). For instance, gang members are more likely to settle in communities that share similar social and demographic traits. This dynamic helps in explaining why activities such as production, trafficking, and governance (Breuer and Varese, 2023; Campana and Varese, 2018), along with related criminal behaviors, are more prevalent in certain urban areas than in others.
@@
The primary policy implication of the geographic concentration and temporal persistence of crime mainly involves the adoption of targeted deterrence strategies, with a particular emphasis on policing. As most crimes are concentrated in small geographic areas (Telep and Weisburd, 2018; Weisburd et al., 2004), law enforcement approaches must be tailored to this reality. One of the most popular strategies is the hot spots approach (e.g., Weisburd and Telep, 2014; Braga et al., 2014), where police resource allocation focuses on geographic areas with a higher number of crimes than the city’s average or where people are at a higher risk of victimization. In this way, police resources could be allocated to just 5% of its block segments to address more than 50% of the total reported incidents it receives (Weisburd, 2015). This approach is mostly guided by predictive techniques that forecast crime in both time and space (Berk, 2021), while the response may involve strategies such as increasing foot patrols, engaging with the local stakeholders, or implementing problem-oriented policing (Braga and Bond, 2008; Taylor et al., 2011; Weisburd et al., 2022).
@@
The main issue with the above perspective is that an over-reliance on purely-predictive methods fails to inform on the cascading effects of hot spot interventions, which include the risk of crime and disorder displacement. Critics of intensive hot spot policing argue that such strategies tend to displace crime rather than effectively reduce it. The literature is inconclusive and cannot rule out positive or negative spillover effects. Braga et al.’s (2019) review shows that studies usually do not detect noticeable negative spillover to other areas. Several analyses even identify benefits that spread from hot spots to adjacent areas (e.g., Weisburd et al., 2006). Both incapacitation and deterrence can explain these results. If the arrested offenders also operated in neighboring areas, the offense is expected to decrease in the hot spots and surrounding areas. On the other hand, some studies document crime displacement when the police strategy focuses on specific areas. For example, a study by Blattman et al. (2017) finds that any crime deterrence in the target area is more than offset by increased crime on the streets within the following 250 meters. In addition, such an analysis unveils heterogeneous effects by crime category. While violent crimes decrease, property crimes increase. The authors conclude that this type of policing pushes property crime around the corner. Similarly, Banerjee et al. (2019) use randomized drunk driving checkpoints to find learning and strategic responses that explain the displacement of crime to other areas.
@@
The lack of a consensus in the aforementioned research studies underscores the pressing need for a more comprehensive understanding of how various crime categories manifest, and interact, in different urban areas across time. Even though crime clustering in time and space is a known phenomenon, an unanimous consensus remains elusive regarding the specific interactions among criminal activities in time, space and across different crime categories. Consequently, an ongoing debate surrounds the tangible benefits of specific intervention policies, particularly those with a pronounced local emphasis (e.g., the hot spots approach) based on black-box predictive methods. In fact, such methods are not designed to provide a model-based and interpretable understanding of complex dependencies in three-way criminal activity tensors. In Section 3, we cover such a gap through a statistical model that can disentangle these dependencies, while preserving a predictive performance competitive with routine-use black-box methods. As such, the proposed perspective can open the avenues for refining the design and evaluation of law enforcement interventions by combining accurate forecasts with a detailed quantitative assessment of the intended or unintended spillover effects that these interventions may trigger. To the best of our knowledge, this important direction has received limited attention in the literature.
@@
As discussed within Section 1, available model-based solutions for the study of criminal activity data (see, e.g., Shoesmith, 2013; Law et al., 2014; Liu and Zhu, 2017; Quick et al., 2018; Clark and Dixon, 2018; Balocchi and Jensen, 2019; Vicente et al., 2023; Balocchi et al., 2023) mostly constrain the analysis to two selected dimensions of the observed three-way tensor. Although such a perspective fails to characterize and learn the full range of dependencies across time, space and crime categories, the available contributions along these directions highlight several fundamental aspects that are the basis of the statistical model we design in Section 3.
More specifically, in line with the discussion in Sections 2.12.2, the design of careful model-based representations for the complex dependencies across time and space is crucial in analyzing criminal activity data. For example, Shoesmith (2013) demonstrates the importance of temporal lags and spatio-temporal interactions when modeling criminal activities. Similarly, Balocchi and Jensen (2019) and Balocchi et al. (2023) show how careful borrowing of information among temporal crime trends across space improves modeling outcomes. Vicente et al. (2023) employ, instead, a penalized spline expansion across locations and times, further allowing for joint modeling of different crime categories. Although such a formulation does not explicitly incorporate graphical dependencies among these different crime categories, it provides relevant support to the importance of devising joint formulations for the entire three-way criminal activity tensor. This support can be found also in the static spatial models by, e.g., Liu and Zhu (2017) and Quick et al. (2018), which highlight how a multivariate perspective focusing on multiple crime categories can unveil structural patterns that a univariate model is unable to detect. Although black-box strategies (see, e.g., Berk, 2021; Mandalapu et al., 2023) can jointly leverage this complex system of dependencies within the observed three-way criminal activity tensor to achieve a remarkable predictive accuracy, the design of a statistical model that can disentangle and infer these dependencies, without sacrificing forecasting performance, is substantially more challenging. Although such a perspective has received limited attention, as discussed in Section 2.2, advancements in this direction are fundamental to devise strategies based not only on accurate forecasts, but also on a careful assessment of the cascading effect of the adopted intervention strategy. The relevance of this endeavor has motivated some efforts towards providing structural interpretation to black-box algorithms in crime analysis, but these interpretations are mostly based on ex-post variable importance measures in the case where external covariates are related to crime activity data under some sort of regression structure (e.g., Zhang et al., 2022).
@@
Following similar practice in the analysis of criminal activity data ( e.g., Shoesmith, 2013; Balocchi and Jensen, 2019), we model the counts $c_{kqt}\in\mathbbm{N}$ of the reported crimes for category $k=1,\dots,K$, within location $q=1,\dots,Q$ at time $t=1,\dots,T$ under the $\log(1+c_{kqt})$ transformation. Let $y_{kqt}$ be the centered target variable obtained by removing from $\log(1+c_{kqt})$ its average over $t$ (keeping $k$ and $q$ fixed), and denote with $\bf{Y}$ the $K\times Q\times T$ criminal activity tensor having entries $y_{kqt}$, for each $k=1,\dots,K$, $q=1,\dots,Q$ and $t=1,\dots,T$. In Section 3.1, we design a statistical model for $\bf{Y}$ which characterizes the temporal dependence among its $K\times Q$ time slices ${\bf Y}_{t}$, $t=1,\dots,T$, via a matrix autoregressive construction that effectively separates space and crime-category effects through a parsimonious, yet interpretable, low-rank bilinear formulation.
The above model, combined with a separable covariance structure for the error term, allows to infer spatial dependence, interlinkages between the different crime categories and persistence over time, without undermining predictive accuracy. Inference for the proposed model proceeds under a Bayesian framework outlined in detail in Section 4.
@@
To devise a model that captures spatial, temporal and crime-category effects within the three-way criminal activity tensor, while facilitating the design of forecasting strategies, a simple building-block option is to rely on a full autoregressive construction for the matrix-valued sequence ${\bf Y}_{t}$, $t=1,\dots,T$. More specifically, let ${\bf y}_{t}=[y_{11t},\dots,y_{kqt},\dots y_{KQt}]^{\intercal}$ denote the vectorized version of the matrix ${\bf Y}_{t}$, for each $t=1,\dots,T$, then it is natural to start by assuming ${\bf y}_{t}=\bm{\Theta}{\bf y}_{t-1}+\bm{\varepsilon}_{t}$, with $\bm{\varepsilon}_{t}\sim\mbox{N}_{KQ}({\bf 0},\bm{\Sigma})$, independently for every $t=1,\dots,T$. In this standard vector autoregressive formulation the lagged values within ${\bf y}_{t-1}$ affect the current ones in ${\bf y}_{t}$ via a full and unstructured $KQ\times KQ$ coefficient matrix $\bm{\Theta}$, which captures the different inter-temporal effects that every combination of location and crime category exerts on each entry in ${\bf y}_{t}$, whereas the $KQ\times KQ$ error covariance matrix $\bm{\Sigma}$ encodes the instantaneous dependencies in reaction to possible shocks.

### v2

@@ Abstract
Criminal activity data are typically available via a three-way tensor encoding the reported frequencies of different crime categories across time and space. The challenges that arise in the design of interpretable, yet realistic, model-based representations of the complex dependencies within and across these three dimensions have led to an increasing adoption of black-box predictive strategies. Although this perspective has proved successful in producing accurate forecasts guiding targeted interventions, the lack of interpretable model-based characterizations of the dependence structures underlying criminal activity tensors prevents from inferring the cascading effects of these interventions across the different dimensions. We address this gap through the design of a low-rank bilinear autoregressive model which achieves comparable predictive performance to black-box strategies, while allowing interpretable inference on the dependence structures of criminal activity reports across crime categories, time, and space. This representation incorporates the time dimension via an autoregressive construction that accounts for spatial effects and dependencies among crime categories through a separable low-rank bilinear formulation. When applied to Chicago police reports, the proposed model showcases remarkable predictive performance and also reveals interpretable dependence structures unveiling fundamental crime dynamics. These results facilitate the design of more refined intervention policies informed by the cascading effects of the policy itself.
@@
The empirical analysis of criminal activities plays an important role in both theory development and public policy considerations. Theoretical models of crime indicate that criminal incidents do not occur randomly, but rather exhibit complex interdependencies across spatial areas, time, and crime categories (e.g., Brantingham and Brantingham, 2016; Weisburd, 2015). As a result, several methods have been proposed to characterize the intricate system of interdependencies within and across these three dimensions. Available solutions for addressing this endeavor range from interpretable model-based representations (e.g., Shoesmith, 2013; Law et al., 2014; Liu and Zhu, 2017; Quick et al., 2018; Clark and Dixon, 2018; Balocchi and Jensen, 2019; Vicente et al., 2023; Balocchi et al., 2023) to more sophisticated black-box predictive strategies (e.g., Guo and Wu, 2012; Chun et al., 2019; Walczak, 2021; Berk, 2021; Zhang et al., 2022; Liang et al., 2022; Mandalapu et al., 2023; Wang et al., 2025). Although statistical models allow, in principle, for interpretable inference on the complex dependence structures underlying the observed criminal activity tensors, the validity of the final conclusions is typically undermined by the challenges that arise in the design of sufficiently flexible, yet interpretable, model-based representations for these dependence structures. As such, the existing statistical models often constrain the analysis by focusing on two dimensions of the three-way tensor, fixing the third at a selected crime category, a single geographic location, or a pre-specified time interval. Although this perspective has led to important advancements, constraining the analysis only to two dimensions fails to characterize the full set of interdependencies underlying the criminal activity tensor, thereby limiting the policy impact of the resulting inferences. Addressing these issues would require extending the available flexible models for two dimensions to similarly effective formulations for the entire three-way tensor. In fact, instead of exploring such a direction, most efforts in the recent years have been devoted to the development and implementation of black-box strategies within a predominantly prediction-oriented framework (see, e.g., Berk, 2021; Mandalapu et al., 2023). These strategies have proved successful in producing accurate forecasts guiding targeted intervention policies, such as hot spot policing (Weisburd and Telep, 2014; Braga et al., 2014), but generally lack a model-based construction, and therefore make it challenging to derive meaningful structural insights on the cascading, positive or negative, spillover effects of specific localized interventions (see, e.g., Weisburd et al., 2006; Blattman et al., 2021). Learning these effects is fundamental to devise improved intervention policies that not only rely on the forecasted criminal activities, but are also guided by the expected cascading effects of the planned policy itself.
@@
Motivated by the above discussion, we design in Section 3 a model-based Bayesian representation of three-way criminal activity tensors that achieves a predictive performance comparable to routinely used black-box strategies, while allowing for interpretable inference on the complex dependence structures within and across crime categories, spatial locations and time intervals. This formulation is guided by qualitative and quantitative theories of crime discussed in Section 2, and relies on a temporal autoregressive construction that incorporates spatial effects and graphical dependencies among different crime categories via a separable low-rank bilinear formulation inspired by the general class of multilinear tensor regression models in Hoff (2015). Besides reducing the dimensionality of the problem, this novel perspective to the analysis of crime activity tensors crucially unveils “averaged” inter-temporal dependencies between crime categories within spatial locations as well as among spatial locations for each crime category. Such a separable structure is also considered for the noise component of the model, thereby further uncovering across-categories and across-locations immediate reactions to shocks, possibly arising from targeted hot spot interventions. As clarified in Section 4, inference under the proposed statistical model is performed through a Bayesian approach leveraging a suitably designed Gibbs sampling algorithm. The advantages of our formulation are illustrated in Section 5 via an application to a three-way criminal activity tensor obtained from the Chicago Police Department. Besides achieving a predictive performance comparable to routinely implemented black-box strategies, the proposed model unveils yet unexplored statistical interdependencies between the crime categories and among spatial locations over time that augment the set of information available to those policymakers seeking to understand and prevent complex patterns in criminal activities. As discussed in the concluding remarks in Section 6, our novel approach can be applied in general to criminal activity tensors from any urban environment for which similar data are available.
@@
Crime is mainly an urban phenomenon. Urban agglomerations typically exhibit higher crime rates compared to smaller towns and rural areas (e.g., Glaeser and Sacerdote, 1999). Although larger cities are often associated with higher economic development, and hence more attractive, this is not the only dimension that explains the distribution of crime across cities. In fact, crime reports are typically sorted unequally over different urban locations, with offenses often concentrated by type in relatively small areas. This spatial heterogeneity, coupled with specific temporal persistences, has led social scientists to focus on local determinants rather than individual drivers.
Three traditional criminology perspectives are useful for understanding the connections among time, space and crime categories. Crime opportunity theories, which focus on the factors shaping space-time crime patterns, particularly the clustering of criminal activities (e.g., Brantingham and Brantingham, 2017; Cohen and Felson, 1979); social control theories, where the structural dimensions of a neighborhood influences residents propensity to exert informal social control (Sampson and Groves, 1989; Shaw and McKay, 1942); and theories that simply highlight how local characteristics (e.g., medical marijuana dispensaries, parks, neighborhood stability and socioeconomic disparities) affect incentives for committing a specific crime within a given area (Contreras, 2017; Contreras and Hipp, 2020; Kubrin and Hipp, 2016). It is important to note that crime clusters are not driven by a single factor in isolation (Hipp and Williams, 2020). Rather, the interplay between opportunity, social controls, and local conditions causes urban areas to experience varying levels of temporal exposure to specific types of crimes and violence.
@@
Although it is challenging to separate time and space when analyzing urban crime patterns, the temporal dimension of crime determinants has been documented to play a crucial role in shaping its spatial distribution, alongside factors related to the offender behavior. Several criminogenic features of a geographic area tend to remain stable over time, contributing to the persistence of specific criminal activities (Johnson, 2008). For example, neighborhood characteristics that either promote or inhibit social control are likely to remain constant (Sampson and Groves, 1989). In addition, the success of previous offenses often motivates criminals to continue operating in the same area, capitalizing on the knowledge acquired about available targets and absence of countermeasures (Cohen and Felson, 1979). The literature on near repeat victimization further suggests that offenders are more likely to target locations where similar crimes have been successfully committed in the past (Johnson and Bowers, 2004; Townsley et al., 2003). However, not all crimes are equally susceptible to temporal persistence and spatial clustering (Pease et al., 1998; Johnson and Bowers, 2004; Ratcliffe and Rengert, 2008; Johnson et al., 2009; Haberman and Ratcliffe, 2012; Amemiya et al., 2020; Chainey, 2021; Chainey and Muggah, 2022). In fact, the clustering of urban crime in time and space is shaped by patterns that are heterogeneous across crime categories. In specific urban areas, certain types of offenses are more prevalent than others, leading to a localized concentration of particular crimes (e.g., property crimes versus violent crimes). These patterns are shaped by the spatial distribution of suitable targets and by the presence of deterrence interventions (Cohen and Felson, 1979), which can favor the occurrence of some crimes over others. For instance, the covid-19 pandemic highlighted the role of crime opportunity in explaining the global reduction in street crime (Nivette et al., 2021; Trajtenberg et al., 2024), while at the same time, the literature documented increments in domestic violence and cybercrime (Piquero et al., 2021; Ravindran and Shah, 2023). Empirical evidence suggests that offenders may have increased the use of violence to compensate for the reduced opportunities brought on by the pandemic, thus reshaping the spatial distribution and clustering of violent and non-violent crimes (Díaz et al., 2022). Another example of crime-type clustering heterogeneity results from the strategic decision-making of organized crime groups when selecting or relocating to new territories (Varese, 2011). These groups typically target urban communities that offer favorable conditions for the development of illegal markets, display socio-economic deprivation, and demonstrate a higher degree of homophily (Campana and Meneghini, 2024; Katz and Schnebly, 2011). For instance, gang members are more likely to settle in communities that share similar social and demographic traits. This dynamic helps in explaining why activities such as production, trafficking, and governance (Breuer and Varese, 2023; Campana and Varese, 2018), along with related criminal behaviors, are more prevalent in certain urban areas than in others.
@@
The primary policy implication of the geographic concentration and temporal persistence of crime mainly involves the adoption of targeted deterrence strategies, with a particular emphasis on policing. As most crimes are concentrated in small geographic areas (see, e.g., Telep and Weisburd, 2018; Weisburd et al., 2004), law enforcement approaches are often tailored to this reality. One of the most popular strategies is the hot spots approach (e.g., Weisburd and Telep, 2014; Braga et al., 2014), where police resource allocation focuses on geographic areas with a higher number of crimes than the city’s average or where the people are at a higher risk of victimization. In this way, police resources could be allocated to just 5% of its block segments to address more than 50% of the total reported incidents it receives (Weisburd, 2015). This approach is mostly guided by predictive techniques that forecast crime in both time and space (Berk, 2021), while the response may involve strategies such as increasing foot patrols, engaging with the local stakeholders, or implementing problem-oriented policing (Braga and Bond, 2008; Taylor et al., 2011; Weisburd et al., 2022).
@@
The main issue with the above perspective is that an over-reliance on purely predictive methods fails to inform on the cascading effects of hot spot interventions, which include the risk of crime and disorder displacement. Critics of intensive hot spot policing argue that such strategies tend to displace crime rather than effectively reduce it. The literature is inconclusive and cannot rule out positive or negative spillover effects. Braga et al.’s (2019) review shows that studies usually do not detect noticeable negative spillover to other areas. Several analyses even identify benefits that spread from hot spots to adjacent areas (e.g., Weisburd et al., 2006). Both incapacitation and deterrence can explain these results. If the arrested offenders also operated in neighboring areas, the offense is expected to decrease in the hot spots and surrounding areas. On the other hand, some studies document crime displacement when the police strategy focuses on specific areas. For example, a study by Blattman et al. (2017) finds that any crime deterrence in the target area is more than offset by increased crime on the streets within the following 250 meters. In addition, such an analysis unveils heterogeneous effects by crime category. While violent crimes decrease, property crimes increase. The authors conclude that this type of policing pushes property crime “around the corner”. Similarly, Banerjee et al. (2019) use randomized drunk driving checkpoints to find strategic responses that explain the displacement of crime to other areas.
@@
The lack of a consensus in the above studies underscores the need for a more detailed understanding of how various crime categories manifest, and interact, in different urban areas across time. Even though crime clustering in time and space is a known phenomenon, an unanimous consensus remains elusive regarding the specific interactions among criminal activities in time, space and across crime categories. Consequently, an ongoing debate surrounds the tangible benefits of specific intervention policies, particularly those with a local emphasis (such as the hot spots approach) based on black-box predictive methods. In fact, these methods are not designed to provide a model-based and interpretable understanding of complex dependencies in three-way criminal activity tensors. In Section 3, we cover this gap through a statistical model that can disentangle these dependencies, while preserving a predictive performance competitive with routine use black-box methods. As such, the proposed perspective can open the avenues for refining the design and evaluation of law enforcement interventions by combining accurate forecasts with a detailed quantitative assessment of the spillover effects that these interventions may trigger. To the best of our knowledge, this direction has received limited attention in the literature.
@@
As discussed in Section 1, available model-based solutions for the study of criminal activity data (e.g., Shoesmith, 2013; Law et al., 2014; Liu and Zhu, 2017; Quick et al., 2018; Clark and Dixon, 2018; Balocchi and Jensen, 2019; Vicente et al., 2023; Balocchi et al., 2023) mostly constrain the analysis to two selected dimensions of the observed three-way tensor. Although such a perspective fails to characterize and learn the full range of dependencies across time, space and crime categories, the available contributions along these directions highlight several important aspects that are the basis of the statistical model we design in Section 3.
More specifically, in line with the discussion in Sections 2.12.2, the design of careful model-based representations for the complex dependencies across time and space is crucial in analyzing criminal activity data. For example, Shoesmith (2013) demonstrates the importance of temporal lags and spatio-temporal interactions when modeling criminal activities. Similarly, Balocchi and Jensen (2019) and Balocchi et al. (2023) show how careful borrowing of information among temporal crime trends across space improves modeling outcomes. Vicente et al. (2023) employ, instead, a penalized spline expansion across locations and times, further allowing for joint modeling of different crime categories. Although such a formulation does not explicitly incorporate graphical dependencies among these different crime categories, it provides relevant support to the importance of devising joint formulations for the entire three-way criminal activity tensor. This support can be found also in the static spatial models by, e.g., Liu and Zhu (2017) and Quick et al. (2018), which highlight how a multivariate perspective focusing on multiple crime categories can unveil structural patterns that a univariate model is unable to detect. Although black-box strategies (e.g., Berk, 2021; Mandalapu et al., 2023) can jointly leverage the complex system of dependencies in the observed three-way criminal activity tensor to enhance predictive accuracy, the design of a statistical model that can disentangle and infer these dependencies, without sacrificing forecasting performance, is substantially more challenging. Although such a perspective has received limited attention, as discussed in Section 2.2, advancements along this direction are necessary to devise strategies based not only on accurate forecasts, but also on a careful assessment of the cascading effect of the adopted intervention strategy. The relevance of this endeavor has motivated some efforts towards providing structural interpretation to black-box algorithms in crime analysis, but these interpretations are based on ex-post variable importance measures in the case where external covariates are related to crime activity via a regression structure (e.g., Zhang et al., 2022).
@@
Following similar practice in the analysis of criminal activity data (e.g., Shoesmith, 2013; Balocchi and Jensen, 2019), we model the counts $c_{kqt}\in\mathbbm{N}$ of the reported crimes for category $k=1,\dots,K$, within location $q=1,\dots,Q$ at time $t=1,\dots,T$ under the $\log(1+c_{kqt})$ transformation. Although modeling the original counts is also a possibility, this option would require Poisson or negative Binomial likelihoods, which, however, have witnessed less extensions and tractable methodological advancements in tensor-valued data settings compared to the classical Gaussian framework employed in practice for $\log(1+c_{kqt})$ (Shoesmith, 2013; Balocchi and Jensen, 2019).
As illustrated in Section 5.1 (see Figure 1), in practical applications the log-transformed counts often exhibit systematic time trends and seasonal components, for several combinations of crime categories and spatial locations. These trends and seasonal effects might obscure the system of stochastic dependencies that are the focus of our analysis of the three-way criminal activity tensor. Following standard practice (see, e.g., Hamilton, 1994; Tsay, 2005), we thus pre-process the data by removing the quadratic trends and the seasonal components displayed by the log-transformed series in Figure 1. This is accomplished by subtracting from $\log(1+c_{kqt})$ the term $\hat{\mu}_{kqt}=\hat{\beta}_{0qk}+\hat{\beta}_{1qk}t+\hat{\beta}_{2qk}t^{2}+\hat{\bm{\gamma}}^{\intercal}_{kq}{\bf d}_{t}$, for every $t=1,\ldots,T$, $k=1,\dots,K$ and $q=1,\dots,Q$, where $\hat{\bm{\gamma}}_{kq}$ contains the monthly seasonal effects, while ${\bf d}_{t}$ is a vector of dummy variables which identify the month associated with the time index $t$. The estimates of these coefficients can be readily obtained via ols applied separately, for every pair $(k,q)$, to the regression model having response $\log(1+c_{kqt})$ and predictor ${\beta}_{0qk}+{\beta}_{1qk}t+{\beta}_{2qk}t^{2}+{\bm{\gamma}}^{\intercal}_{kq}{\bf d}_{t}$, for $t=1,\ldots,T$.
It is important to emphasize that the above de-trending and seasonal adjustment does not prevent from predicting crime levels on the original scale. In fact, as clarified in Section 5.2.1, it suffices to add $\hat{\mu}_{kqt}$ to the forecasts of the statistical model designed in Section 3.1 for the de-trended and seasonally-adjusted log-transformed counts, which we denote as $y_{kqt}$. More specifically, let $\bf{Y}$ be the $K\times Q\times T$ criminal activity tensor with entries $y_{kqt}$, for each $k=1,\dots,K$, $q=1,\dots,Q$ and $t=1,\dots,T$. In Section 3.1, we design a statistical model for $\bf{Y}$ that characterizes the temporal dependence among its $K\times Q$ (crime-category – location) slices ${\bf Y}_{t}$, at time $t=1,\dots,T$, via a matrix autoregressive construction that separates space and crime-category effects through a parsimonious, yet interpretable, low-rank bilinear formulation. This model, combined with a separable covariance structure for the error term, allows to infer spatial dependence, interlinkages between the different crime categories and persistence over time, without undermining predictive accuracy. Inference for the proposed model proceeds under a Bayesian framework outlined in Section 4.
@@
To devise a model that captures spatial, temporal and crime-category effects within the three-way criminal activity tensor, while facilitating the design of forecasting strategies, a simple building-block option is to rely on a full autoregressive construction for the matrix-valued sequence ${\bf Y}_{t}$, $t=1,\dots,T$. More specifically, let ${\bf y}_{t}=\mbox{vec}({\bf Y}_{t})=[y_{11t},\dots,y_{kqt},\dots y_{KQt}]^{\intercal}$ be the vectorized version of the matrix ${\bf Y}_{t}$, for each $t=1,\dots,T$, then it is natural to start by assuming ${\bf y}_{t}=\bm{\Theta}{\bf y}_{t-1}+\bm{\varepsilon}_{t}$, with $\bm{\varepsilon}_{t}\sim\mbox{N}_{KQ}({\bf 0},\bm{\Sigma})$, independently for every $t=1,\dots,T$. In this standard vector autoregressive formulation the lagged values within ${\bf y}_{t-1}$ affect the current ones in ${\bf y}_{t}$ via a full and unstructured $KQ\times KQ$ coefficient matrix $\bm{\Theta}$, which captures the different inter-temporal effects that every combination of location and crime category exerts on each entry in ${\bf y}_{t}$, whereas the $KQ\times KQ$ error covariance matrix $\bm{\Sigma}$ encodes the instantaneous dependencies in reaction to possible shocks.